Find an equation of the line passing through
the point (-6, 5) with slope m = 5.

Answer by DrBeeee(684)

The slope-intercept form of a straight line is
(1) y = m*x + b
You are given m=5 so (1) becomes
(2) y = 5*x + b
You are given that this line passes through the point (-6,5) so (2) becomes
(3) 5 = 5*(-6) + b or
(4) 5 = -30 + b or
(5) b = 5 + 30 or
(6) b = 35, therefore the required straight line is
(7) y = 5*x + 35
Let's check this with the point (-6,5).
Is (5 = 5*(-6) + 35)?
Is (5 = -30 + 35)?
Is (5 = 5)? Yes
Answer: That equation of the line with slope 5, and passes through (-6,5) is y = 5x + 35.